# Focused Angular $N$-Body Event Generator (FANG)

### v1

@@ IV.1 Full phase space calculation
Testing the full phase-space calculation verifies that the $M$-generation is correctly implemented in the FANG algorithm. This is done by comparing full phase-space volumes computed with FANG (see Eq. 21) to those obtained with the widely used event generator RAMBO [2]. The results summarized in Table 1 show excellent agreement between the two codes. In most cases, RAMBO is slightly more efficient than FANG, although at low energies FNAG becomes comparatively more competitive.
@@
Table 1: Phase-space calculations for $P(0,0,5,13)$ decaying into final states with the specified masses. The HAZMA [16] results, which are an implementation of RAMBO, have been multiplied by $(2\pi)^{3n-4}$.
Masses FANG HAZMA
$10^{8}$ events $10^{8}$ events
1,1,1 $1.469\pm 0.015 141.4$$1.458\pm 0.002$
1,1,1,1 $3285.76\pm 0.36$ $3286.81\pm 0.10$
1,1,1,1,1 $26628.1\pm 3.0$ $26630.0\pm 1.2$
1,1,1,1,1,1 $80467.7\pm 9.4$ $80499.3\pm 5.0$
2,2,2,2,2,1.9 $4.87011E-7\pm 5.6E-11$ $4.87115\times 10^{-7}\pm 5.8\times 10^{-11}$
IV.2 Partial phase space event generation
@@
To test the partial phase-space event generation, the angular and energy distributions of the outgoing particles are compared between FANG, which generates only events within the designated solid angles, and a full phase-space Monte Carlo, where cuts are applied to select events in which the restricted particles fall within the designated solid angles. The full phase-space Monte Carlo used for this comparison is GENBOD, as implemented in the TGenPhaseSpace class of the ROOT Data Analysis Framework [17]. The test scenario considers a parent particle with four-momentum $P(P_{x},P_{y},P_{z},E)=(0,0,5,13)$ decaying into five particles, each of mass $m_{i}=1$, at the center of the laboratory coordinate system. Three of the particles are constrained. The first, with momentum $p_{1}$, is required to reach a circular “detector” of radius 0.2 centered along the $z$ axis at the coordinates $(0,0,0.5)$. The second, with momentum $p_{2}$, is required to reach a circular “detector” of radius 0.3 centered along the $x$ axis at the coordinates $(0.5,0,0)$. The third particle is constrained to lie within a solid-angle strip defined by $0.75<\cos\theta<0.75$ and $0.1\pi<\varphi<0.9\pi$. The phase space of the unconstrained full Monte Carlo is given in Table 1 as $26628.1\pm 3.0$, while the phase space of the constrained Monte Carlo, calculated using Eq. 35, is $4.4544\pm 0.0061$, i.e., about $\sim 5000$ times smaller. This implies that achieving the same statistical precision would require running the full Monte Carlo with roughly $5000$ times more events than FANG. For this comparison, a relatively large solid angle was deliberately chosen; however, in some scenarios the constrained phase space can be as small as $10^{-9}$ of the full phase space, in which case relying on a full Monte Carlo with cuts becomes impractical.
@@
Figure 2 shows perfect agreement between the $\cos\theta$ distributions of all five particles obtained from the full simulation with cuts and from the FANG algorithm, which generates only events within the designated solid angles. For the tests presented here, FANG was run with $10^{6}$ events, while the full GENBOD simulation required $10^{9}$ events and still yielded five times fewer statistics after cuts. Similar levels of agreement are observed in the $\varphi$ distributions (Fig. 3) and in the energy distributions (Fig. 4).
@@
Figure 4: Comparison of the energy distributions for all five particles between the full simulation with cuts and the FANG algorithm, which generates only events within the designated solid angles.
IV.3 Differential cross section calculation
Having established the correctness of the total phase-space integration, the next step is to verify that the method also reproduces physical observables by integrating the phase space with the squared matrix amplitude, as required for the calculation of differential cross sections. The differential cross-section calculation is tested using a textbook example of $e^{-}\mu^{-}\rightarrow e^{-}\mu^{-}$ scattering taken from Quarks & Leptons by Halzen and Martin [18] (see Fig. 5). The differential cross section for this process can be derived analytically and is therefore well suited for comparison with calculations performed using the FANG algorithm.
@@ IV.3 Differential cross section calculation
The spin-averaged $e^{-}\mu^{-}\longrightarrow e^{-}\mu^{-}$ squared amplitude for this process, as given in Ref. [18], is calculated to be
@@
Here $E$ and $E^{\prime}$ denote the energies of the incoming and outgoing electrons, respectively, and the laboratory frame is defined as the frame in which the muon is initially at rest. Figures 6 and 7 compare the analytic expression of Eq. 30 (red line) with the FANG calculation (black points). The Monte Carlo calculation was performed using the point-generation option, in which events are generated with the electron directed to a specific point in each event, and by inserting $|\mathcal{M}|^{2}$ from Eq. 29 into Eq. 26. For high energies, such as a $3~\text{GeV}$ electron (Fig. 6), the analytical and FANG calculations coincide. At lower energies, such as a $3~\text{MeV}$ electron (Fig. 7), a divergence appears, as expected, since the analytical calculation neglects all terms proportional to the electron mass, while the Monte Carlo includes the full expression. A closely related example is elastic $ep\rightarrow ep$ scattering taken from [19],
@@
with $\mu=2.793$ the proton magnetic moment. Based on the above definitions, the FANG calculation of the differential cross section at selected kinematic points can be benchmarked against the Rosenbluth formula, providing a direct test of the method. The results of this comparison are shown in Fig. 8. All three figures also show excellent agreement with the angular distributions of the generated events, displayed as histograms in Figs. 68. The events were generated using the user-defined option, which allows the constraining vectors to be chosen anisotropically in the laboratory frame, provided they are subsequently reweighted to ensure isotropy. In this example, the vectors were generated with an angular distribution proportional to $1/\sin^{4}(\tfrac{\theta}{2})$. The events were then reweighted by a factor of $\sin^{4}(\tfrac{\theta}{2})$, which cancels the corresponding factor in $|\mathcal{M}|^{2}$. In this way, the variation of the event weights due to the strong angular dependence of the squared amplitude is removed, significantly reducing the statistical uncertainty in the weighted sum.
Figure 6: Differential $e^{-}\mu^{-}$ elastic cross-section, showing agreement between the analytic expression of Eq. 30 (red line) and the FANG calculation (black points). The black points are obtained by inserting $|\mathcal{M}|^{2}$ from Eq. 29 into Eq. 26 and using event point generation.
Figure 7: Expected divergence in the differential $e^{-}\mu^{-}$ elastic cross section between the analytic expression of Eq. 30 (red line), which neglects the $m_{e}$ terms, and the FANG calculation (black points), obtained by inserting $|\mathcal{M}|^{2}$ from Eq. 29 with all $m_{e}$ terms included into Eq. 26.
Figure 8: Differential $e^{-}p$ elastic cross section, showing agreement between the Rosenbluth formula (Eq. 8.17 in [18], red line) and the FANG calculation (black points). The black points are obtained by inserting $|\mathcal{M}|^{2}$ into Eq. 26 and using event point generation.
IV.4 Comparison with Lepton–Nucleus Scattering Models
GENIE [20] is a widely used neutrino event generator that simulates interactions of neutrinos with all nuclear targets and flavors across a broad energy range, from approximately $1$ MeV up to $1$ PeV. In GENIE, the relativistic Fermi gas (RFG) nuclear model is applied to all processes. The most stringent tests of the RFG model come from electron scattering experiments [21].
Within the GENIE framework, the electron energy transfer distribution in the lab frame, $\nu=E_{e}-E^{\prime}_{e}$, in scattering at a fixed (large) angle is strongly correlated with the nucleon momentum distribution described by the RFG model. To obtain sufficient statistics from GENIE simulations restricted to, for example, electron scattering at $60\pm 0.5^{\circ}$, it is necessary to impose a $Q^{2}$ cut. However, this cut in turn limits the maximum accessible energy transfer, preventing simulation of the full measured distribution. Moreover, after applying a $Q^{2}$ cut, the majority of simulated events correspond to forward-going electrons that do not contribute to the desired kinematics. As a result, the process becomes both inefficient and computationally expensive.
@@ IV.4 Comparison with Lepton–Nucleus Scattering Models
IV.4.1 Implementing the GENIE Physics Model in FANG
@@
Each channel has its own cross section. Extracting the exact squared matrix element $|\mathcal{M}|^{2}$ that reproduces the GENIE cross section for each channel is beyond the scope of this simple demonstration. Instead, the total number of events in FANG was normalized to the GENIE sample for each channel separately. The following simplified amplitudes were used: for the QE and RES channels, the same elastic scattering amplitude as Eq. 31 was employed with the appropriate four-momenta; the MEC channel was generated using isotropic phase space ($|\mathcal{M}|^{2}=1$); and for the DIS channel, the amplitude for $ep\rightarrow eX$ scattering was taken from Ref. [18]:
@@
where $p$ is the four-momentum of the initial nucleon, $p^{\prime}$ the four-momentum of the outgoing hadronic system, and $M$ its mass. For simplicity, the following relations are used: the structure functions are written as $\nu W_{2}=F_{2}$ and $MW_{1}=\tfrac{F_{2}}{2x}$, the Bjorken scaling variable is defined as $x=\tfrac{Q^{2}}{2M\nu}$, and the neutron structure function is approximated by $F^{en}_{2}=F^{ep}_{2}(1-0.75x)$. where the proton structure function $F^{ep}_{2}(x,Q^{2})$ is taken from the ALLM97 parametrization [22, 23], a 23-parameter fit to world data.
IV.4.2 Comparison of FANG and GENIE Energy Transfer Results
@@
Figure 10 shows further comparisons at different scattering angles. At $\theta=37.5\pm 0.5^{\circ}$ the advantage of FANG namely, the ability to generate the full energy transfer distribution without artificial cutoffs from a $Q^{2}$ requirement is clearly demonstrated.
@@
Generating $10^{7}$ events in this setup with FANG requires only 2–3 minutes on a single CPU. For realistic physics analyses, however, the amplitudes should be replaced with those corresponding to the exact cross-section calculations implemented in GENIE, rather than the simplified approximations used here.
V Summary
@@
By combining exact kinematics with computational efficiency, FANG offers a versatile tool for differential cross-section and decay-rate calculations in high-energy and nuclear physics. Its robustness makes it well suited for applications ranging from precision tests of the Standard Model to detector design and analysis in particle and nuclear experiments.
@@ Acknowledgements.
We thank Adi Ashkenazi and her research group at Tel Aviv University for their generous help and insightful discussions on the GENIE-related aspects of this work.
Appendix A Integration over Detector geometry
For partial phase-space event generation, where a constrained particle is restricted to a finite solid angle $\Delta\Omega_{k}$ in the laboratory frame, a new constraining vector $\hat{r}_{k}$ is generated isotropically within $\Delta\Omega_{k}$ for each event. In these cases, the event phase space restricted to finite solid angles is given by the angle-integrated form of Eq. 21.
For differential calculations, the algorithm supports the following configurations:

### v2

@@ 4.1 Full phase space calculation
Testing the full phase-space calculation verifies that the $M$-generation is correctly implemented in the FANG algorithm. This is done by comparing full phase-space volumes computed with FANG (see Eq. 13) to those obtained with HAZMA  coogan_hazma_2020 , which is an implementation of the widely used event generator RAMBO kleiss_new_1986 . The results summarized in Table 1 show excellent agreement between the two codes. In most cases, RAMBO is slightly more efficient than FANG, although at the non-relativistic limit, FANG becomes comparatively more competitive.
@@
Table 1: Phase-space volume calculations for $P(P_{x},P_{y},P_{z},E)=(0,0,5,13)$ (same units as mass), decaying into final states with the specified masses. The HAZMA results have been multiplied by $(2\pi)^{3n-4}$.
Number Masses FANG HAZMA
of bodies values $10^{8}$ events $10^{8}$ events
3 1,1,1 $1.469\pm 0.015 141.4$$1.4581\pm 0.0020$
4 1,1,1,1 $3285.76\pm 0.36$ $3286.81\pm 0.10$
5 1,1,1,1,1 $26628.1\pm 3.0$ $26630.0\pm 1.2$
6 1,1,1,1,1,1 $80467.7\pm 9.4$ $80499.3\pm 5.0$
6 2,2,2,2,2,1.9 $4.87011\times 10^{-7}\pm 5.6\times 10^{-11}$ $4.87115\times 10^{-7}\pm 5.8\times 10^{-11}$
4.2 Partial phase space event generation
@@
To test the partial phase-space event generation, the angular and energy distributions of the outgoing particles are compared between FANG, which generates only events within the designated solid angles, and a full phase-space Monte Carlo, where cuts are applied to select events in which the restricted particles fall within the designated solid angles. The full phase-space Monte Carlo used for this comparison is GENBOD, as implemented in the TGenPhaseSpace class of the ROOT Data Analysis Framework canal_root-projectroot_2025 . The test scenario considers a parent particle with four-momentum $P(P_{x},P_{y},P_{z},E)=(0,0,5,13)$ decaying into five particles, each of mass $m_{i}=1$, at the origin of the laboratory coordinate system. Three of the particles are constrained. The first, with momentum $p_{1}$, is required to reach a circular “detector” of radius 0.2 centered along the $z$ axis at the coordinates $(0,0,0.5)$ (same units of length as the radius). The second, with momentum $p_{2}$, is required to reach a circular “detector” of radius 0.3 centered along the $x$ axis at the coordinates $(0.5,0,0)$. The third particle is constrained to lie within a solid-angle strip defined by $0.75<\cos\theta<0.75$ and $0.1\pi<\varphi<0.9\pi$. The phase space of the unconstrained full Monte Carlo is given in Table 1 as $26628.1\pm 3.0$, while the phase space of the constrained Monte Carlo, calculated using Eq. 25, is $4.4544\pm 0.0061$, i.e., about $\sim 5000$ times smaller. This implies that achieving the same statistical precision would require running the full Monte Carlo with roughly $5000$ times more events than FANG. For this comparison, a relatively large solid angle was deliberately chosen; however, in some scenarios the constrained phase space can be as small as $10^{-9}$ of the full phase space, in which case relying on a full Monte Carlo with cuts becomes impractical.
@@
Fig. 2 shows perfect agreement between the $\cos\theta$ distributions of all five particles obtained from the full simulation with cuts and from the FANG algorithm, which generates only events within the designated solid angles. For the tests presented here, FANG was run with $10^{6}$ events, while the full GENBOD simulation required $10^{9}$ events and still yielded five times fewer statistics after cuts. Similar levels of agreement are observed in the $\varphi$ distributions (Fig. 3) and in the energy distributions (Fig. 4).
@@
Figure 4: Comparison of the energy distributions (same units as mass) for all five particles between the full simulation with cuts and the FANG algorithm, which generates only events within the designated solid angles.
4.3 Differential cross section calculation
Having established the correctness of the total and partial phase-space integration, the next step is to verify that the method also reproduces physical observables by integrating the phase space with the squared matrix amplitude, as required for the calculation of differential cross sections. The differential cross section calculation is tested using a textbook example of $e^{-}\mu^{-}\rightarrow e^{-}\mu^{-}$ scattering taken from Quarks & Leptons by Halzen and Martin halzen_quark_2008 (see Fig. 5). The differential cross section for this process can be derived analytically and is therefore well suited for comparison with calculations performed using the FANG algorithm.
@@ 4.3 Differential cross section calculation
The spin-averaged $e^{-}\mu^{-}\longrightarrow e^{-}\mu^{-}$ squared amplitude for this process, as given in Ref. halzen_quark_2008 , is calculated to be
@@
Here $E$ and $E^{\prime}$ denote the energies of the incoming and outgoing electrons, respectively, and the laboratory frame is defined as the frame in which the muon is initially at rest.
Fig. 6 and Fig. 7 compare the analytic expression of Eq. 31 (red line) with the FANG differential cross section calculation using point generation routine at specific points (black points), and the blue histogram of events generated by FANG using the user-defined routine. The user-defined routine allows the constraining vectors to be chosen anisotropically in the laboratory frame, provided the events are subsequently reweighted to ensure isotropy. In this example, the vectors were generated with an angular distribution proportional to $1/\sin^{4}(\tfrac{\theta}{2})$. The events were then reweighted by a factor of $\sin^{4}(\tfrac{\theta}{2})$, which cancels the corresponding factor in $|\mathcal{M}|^{2}$. In this way, the variation of the event weights due to the strong angular dependence of the squared amplitude is removed, significantly reducing the statistical uncertainty in the weighted sum. In all cases the calculation was done by inserting $|\mathcal{M}|^{2}$ from Eq. 30 into Eq. 27. For high energies, such as a $3~\text{GeV}$ electron (Fig. 6), the analytical and FANG calculations coincide. At lower energies, such as a $3~\text{MeV}$ electron (Fig. 7), a divergence appears, as expected, since the analytical calculation neglects all terms proportional to the electron mass, while the Monte Carlo includes the full expression.
Another closely related example is elastic $ep\rightarrow ep$ scattering taken from eichmann_qcd_2020 ,
@@
with $\mu=2.793$ the proton magnetic moment. Based on the above definitions, the FANG calculation of the differential cross section at selected kinematic points can be benchmarked against the Rosenbluth formula rosenbluth_high_1950 , providing a direct test of the method. The results of this comparison are shown in Fig. 8. All three figures also show excellent agreement with the angular distributions of the generated events, displayed as histograms in Figs. 68.
Figure 6: Differential $e^{-}\mu^{-}$ elastic cross section, showing agreement between the analytic expression of Eq. 31 (red line), the FANG calculation (black points) and the (blue histogram) of events generated by FANG.
Figure 7: Expected divergence in the differential $e^{-}\mu^{-}$ elastic cross section between the analytic expression of Eq. 31 (red line), which neglects the $m_{e}$ terms, and the FANG calculation (black points), and the (blue histogram) of events generated by FANG.
Figure 8: Differential $e^{-}p$ elastic cross section, showing agreement between the Rosenbluth formula (Eq. 8.17 in halzen_quark_2008 , red line) and the FANG calculation (black points), and the (blue histogram) of events generated by FANG.
4.4 Testing FANG with Lepton–Nucleus Scattering Models
GENIE andreopoulos_genie_2010 is a widely used event generator for interactions of neutrinos (or electrons) with all nuclear targets and flavors across a broad energy range, from approximately $1$ MeV up to $1$ PeV. In GENIE, the relativistic Fermi gas (RFG) nuclear model is applied to all processes. The most stringent tests of the RFG model come from electron scattering experiments papadopoulou_inclusive_2021 .
Within the GENIE framework, the electron energy transfer distribution in the laboratory frame, $\nu=E_{e}-E^{\prime}_{e}$, in scattering at a fixed (large) angle is strongly correlated with the nucleon momentum distribution described by the RFG model. To obtain sufficient statistics from GENIE simulations restricted to, for example, electron scattering at $60\pm 0.5^{\circ}$, it is necessary to impose a $Q^{2}$ cut. However, this cut in turn limits the maximum accessible energy transfer, preventing simulation of the full measured distribution. Moreover, after applying a $Q^{2}$ cut, the majority of simulated events correspond to forward-going electrons that do not contribute to the desired kinematics. As a result, the process becomes both inefficient and computationally expensive.
@@ 4.4 Testing FANG with Lepton–Nucleus Scattering Models
4.4.1 Implementing the GENIE Physics Model in FANG
@@
Each channel has its own cross section. Extracting the exact squared matrix element $|\mathcal{M}|^{2}$ that reproduces the GENIE cross section for each channel is beyond the scope of this simple demonstration. Instead, the total number of events in FANG was normalized to the GENIE sample for each channel separately. The following simplified amplitudes were used: for the QE and RES channels, the same elastic scattering amplitude as Eq. 32 was employed with the appropriate four-momenta; the MEC channel was generated using isotropic phase space ($|\mathcal{M}|^{2}=1$); and for the DIS channel, the amplitude for $ep\rightarrow eX$ scattering was taken from Ref. halzen_quark_2008 :
@@
where $p$ is the four-momentum of the initial nucleon, $p^{\prime}$ the four-momentum of the outgoing hadronic system, and $M$ its mass. For simplicity, the following relations are used: the structure functions are written as $\nu W_{2}=F_{2}$ and $MW_{1}=\tfrac{F_{2}}{2x}$, the Bjorken scaling variable is defined as $x=\tfrac{Q^{2}}{2M\nu}$, and the neutron structure function is approximated by $F^{en}_{2}=F^{ep}_{2}(1-0.75x)$. where the proton structure function $F^{ep}_{2}(x,Q^{2})$ is taken from the ALLM97 parametrization abramowicz_allm_2004 ; abramowicz_parametrization_1991 , a 23-parameter fit to world data.
4.4.2 Comparison of FANG and GENIE Energy Transfer Results
@@
Fig. 10 shows further comparisons at different scattering angles. At $\theta=37.5\pm 0.5^{\circ}$ the advantage of FANG namely, the ability to generate the full energy transfer distribution without artificial cutoffs from a $Q^{2}$ requirement is clearly demonstrated.
@@
For realistic physics analyses, however, the matrix element should be replaced with those corresponding to the exact cross section calculations implemented in GENIE, rather than the simplified approximations used here.
5 Summary
@@
By combining exact kinematics with computational efficiency, FANG offers a versatile tool for differential cross section and decay-rate calculations in high-energy and nuclear physics. Its robustness makes it well suited for applications ranging from precision tests of the Standard Model to detector design and analysis in particle and nuclear experiments.
@@ Acknowledgements.
We thank Adi Ashkenazi and her research group at Tel Aviv University for their generous help and insightful discussions on the GENIE-related aspects of this work. We also thank Israel Mardor for helpful comments on an earlier version of this manuscript